List of Figures

1 Regular figure 1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 2 Continued figure 2A . . . . . . . . . . . . . . . . . . . . . . . . . 5 3 This is a simple figure. . . . . . . . . . . . . . . . . . . . . . . . . 14

“numberline3(a) . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 “numberline3(b) . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 “numberline3(c) . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 “numberline3(d) . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

4 This is a continued figure. . . . . . . . . . . . . . . . . . . . . . . 14

List of Tables

1 Regular table 1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 2 Continued table 2A . . . . . . . . . . . . . . . . . . . . . . . . . . 10 3 This is a simple table. . . . . . . . . . . . . . . . . . . . . . . . . 15

“numberline4(a) . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 “numberline4(b) . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 “numberline4(c) . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 “numberline4(d) . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 “numberline4(e) . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 “numberline4(f) . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 “numberline4(g) . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 “numberline4(h) . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 “numberline4(i) . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 “numberline4(j) . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 “numberline4(k) . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 “numberline4(l) . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

4 This is a continued table. . . . . . . . . . . . . . . . . . . . . . . 15 “numberline4(a) . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 “numberline4(b) . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 “numberline4(c) . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 “numberline4(d) . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

First we test the regular figures. The first figure is Figure 1 on page 4. The second is Figure 2 and appears on pages 5, 6, 7, and 8.

Then we test the regular tables. The first table is Table 1 on page 9. The second is Table 2 and appears on pages 10, 11, 12, and 13.

Next we look at the subfigures in a simple figure. They are Figures 3(a), 3(b), 3(c), and 3(d), which are part of Figure 3 on page 14.

The continued subfigures are contained in three parts. The first is Figure 4 on page 14 which contains the Subfigures 4(a), 4(b), 4(c), and 4(d). The second is Figure 4 on page 14 which contains the Subfigures 4(e), 4(f), 4(g), and 4(h). The third is Figure 4 on page 15 which contains the Subfigures 4(i), 4(j), 4(k), and 4(l).

Next we look at the subtables in a simple table environment. They are Tables 4(a), 4(b), 4(c), and 4(d), which are part of Table 3 on page 15.

The continued subtables are contained in three parts. The first is Table 4 on page 15 which contains the Subtables 5(a), 5(b), 5(c), and 5(d). The second is Table 4 on page 16 which contains the Subtables 5(e), 5(f), 5(g), and 5(h). The third is Table 4 on page 16 which contains the Subtables 5(i), 5(j), 5(k), and 5(l).

Regular figure 1 Figure 1: Regular figure 1

Continued figure 2A Figure 2: Continued figure 2A

Continued figure 2B Figure 2: Continued figure 2B

Continued figure 2C Figure 2: Continued figure 2C

Continued figure 2D Figure 2: Continued figure 2D Table 1: Regular table 1

Regular table 1 Table 2: Continued table 2A

Continued table 2A Table 2: Continued table 2B

Continued table 2B Table 2: Continued table 2C

Continued table 2C Table 2: Continued table 2D

Continued table 2D

Subfigure 1A_{(a)}  Subfigure 1B_{(b)}
Subfigure 1C_{(c)}  Subfigure 1D_{(d)}

Figure 3: This is a simple figure.

Subfigure 2A_{(a)}  Subfigure 2B_{(b)}
Subfigure 2C_{(c)}  Subfigure 2D_{(d)}

Figure 4: This is a continued figure.

Subfigure 2E_{(e)}  Subfigure 2F_{(f)}
Subfigure 2G_{(g)}  Subfigure 2H_{(h)}

Figure 4: This is a continued figure (cont.)

Subfigure 2I_{(i)}  Subfigure 2J_{(j)}
Subfigure 2K_{(k)}  Subfigure 2L_{(l)}

Figure 4: This is a continued figure (cont.) Table 3: This is a simple table.

Subtable 1A_{(a)}  Subtable 1B_{(b)}
Subtable 1C_{(c)}  Subtable 1D_{(d)}

Table 4: This is a continued table.

Subtable 2A_{(a)}  Subtable 2B_{(b)}
Subtable 2C_{(c)}  Subtable 2D_{(d)}

Table 4: This is a continued table (cont.)

Subtable 2E_{(e)}  Subtable 2F_{(f)}

Subtable 2G_{(g)}  Subtable 2H_{(h)}

Table 4: This is a continued table (cont.)

Subtable 2I_{(i)}  Subtable 2J_{(j)}

Subtable 2K_{(k)}  Subtable 2L_{(l)}